Problem 2 An experiment was conducted to determine the affects of temperature on the mean yield of a chemical process. The data from the experiment (in the order collected) are shown in the following table Temperature (F) 300 550 500 300 250 375 300 250 500 550 375 500 450 375 550 450 250 450 Yield (mg) 25.0 36.3 25.9 21.8 29.4 19.6 33.2 25.1 26.2 38.3 12.7 39.1 21.8 17.5 35.1 26.3 30.5 19.3 a) Does temperature affect the mean yield? Conduct an ANOVA. Use α-0.05 to determine significance. Be sure to analyze the residuals from your ANOVA b) Which temperature(s) produced the statistically significant maximum mean yield? Use a-0.05 c) Fit an appropriate regression model relating temperature to vield. Be sure to analyze the residuals from your regression analysis d) Predict the yield for a temperature of 395 degrees e) Construct a 95% prediction interval (two tailed) for a temperature of 395 degrees
Answer #1

a) Apply One way Anova on Excel.

Group 1 = Yields at 300 Degree

Group 2 = Yields at 550 Degree

Group 3 = Yields at 500 Degree

Group 4 = Yields at 250 Degree

Group 5 = Yields at 375 Degree

Analysis of Variance Results

F-statistic value = 6.60854

P-value = 0.00356

Null Hypothesis:- Temperature does not affects yield.

Alternate Hypothesis:- Temperature affects yield

SInce P- value < 0.05, hence Null Hypothesis is true, which means Temperature does not affects yield

Data Summary Groups Group 1 Group 2 Group 3 Group 4 Group 5 Group 6 Mean 26.6667 36.5667 30.4 28.3333 16.6 22.4667 Std. Dev 5.8799 1.6166 7.5359 2.8537 3.5369 3.5473 Std. Error 3.3948 0.9333 4.3509 1.6476 2.0421 2.048 2 2
b)

The above chart shows Mean is highest for group 2, hence Yield is maximum at 550 degree

c)Step 1: Find X- Y and X2 as it was done in the table below 300 300 300 550 550 550 500 500 500 250 250 250 375 375 375 450 450 450 25 21.8 33.2 36.3 38.3 35.1 25.9 26.2 39.1 29.4 25.1 30.5 19.6 12.7 17.5 21.8 26.3 19.3 7500 6540 9960 19965 21065 19305 12950 13100 19550 7350 6275 7625 7350 4762.5 6562.5 9810 11835 8685 90000 90000 90000 302500 302500 302500 250000 250000 250000 62500 62500 62500 140625 140625 140625 202500 202500 202500

Step 2: Find the sum of every column: x = 7275 , Y 483.1 , ух.ץ = 200190 , ух-3144375 Step 3: Use the following equations to find a and b: ΣΥΣΧ2-ΣΧΣΧΥ 483.1-3144375-7275-200190 p17.06 18-3144375-72752-一~: 200190 17.06 a= n . ΣΧΥ-Σχ . ΣΥ 18-200190-7275-483.1 n.ΣΧ2-(ΣΧ)2 Step 4: Substitute a and b in regression equation formula 18-3144375-(7275)2 y = 17.064 0.024-

y 17.06 + 0.024. X    is the equation.

d) for x= 395 degrees

y = 17.06+0.024*(295) = 24.14mg

e)